Project

General

Profile

NDOS Commissioning Meeting 2011/01/10

Attendance: Jon, Gary, Kurt, Andrew, Ron, Genaidy, Qiming, Denis, Mat, Athans, Susan, Minerba, Rick, Neil, Greg, Leon, Sue, Chad, Peter, Alena

Agenda:
  1. DAQ/Electronics Update & Plans
  2. DCS Status & Plans
  3. DAQ/Electronics Update & Plans
  4. Shifters' report
  5. AOB

DAQ:

  • Peter:
    • latest version of the DAQ software was installed on Fri.:
      • there was some confusion over the HEAD version of DCMProgUtils; a change in a kernel module broke DSO mode.
      • however, Ron claims that this change (additional 1 ms between command and issue of synch pulse) should not have any impact on data taking.
      • need to follow up with Nathan
    • updates to software included improvements to RC and MsgViewer
  • Andrew:
    • DAQ was left unable to take DSO data over the weekend
    • he backed out the change to the kernel module made by Qiming and things seem to be working now (DSO run in progress during the meeting)
  • Peter:
    • tested use of memory map in DCMApp. This seems to be working fine and is in use now. Syst. CPU usage went down from 40% to 20% for DCMs that were hovering around the 40% usage mark. Those that were lower, eg at 20%, stayed at 20%.
    • will attempt to upgrade TDU firmware @ NDOS from version 1.5 to 1.7 today. This should make synching more reliable. However, updating the firmware is non-trivial, as network loading doesn't seem to work. Therefore we need a physical connection to the TDU at NDOS, as well as special equipment (specific laptop + connection device).
    • created a DSO release. If the development version of the SW fails in DSO mode, we can fall back to the DSO release. This requires DAQ expert, however.
  • Sue: noted that over the weekend one DCM suddenly stopped listening to RC heartbeats. Not clear why the run continued ... Jon will investigate.
  • Peter:
    • MB triggers:
      • Denis got spilltimes from MB and compared them to NOvA spilltimes. The spilltimes agree to within a couple of microseconds.
      • Mark still has not found any clear signal of MB events in the data. However, he estimated the number of events we should be expecting and came to the conclusion that the ratio of NuMI to MB events is ~8:1. That's about 1 event/2 days of operations. Also suggested that we get MB spilltimes directly from the accelerator (if possible).
  • Mat:
    • yesterday, 10 APDs were removed:
      • 2 from the top of block 0; visual inspection indicates these are "generally dirty"
      • 8 from the side of block 2; 5 were found to be oily, and these were near where the oil spilled during filling tests; 3 were further upstream, had no oil, but one does have fiber imprints on the face of the APD.
    • it's clear that there is a problem with the optical fibers touching the face of the APDs. How to secure the fibers or keep them from touching the APD?
    • also noted that when fibers are touching the APD, the imprints appear along an edge of the APD. Possibly due to misalignment of the APD, eg put in at an angle?
  • Leon: will look into whether or not we can reliably measure the thickness of the glue which could be the reason fibers are touching APDs.
  • Mat: also looking for ideas of a better method of taking pictures of the APDs and optical fibers.
  • Peter: status of firmware?
  • Ron: working on the switch between external and internal command links.
  • Rick:
    • made a change so that we never send a "," at the wrong time
    • having trouble reconnecting to the clock from the data stream
    • should be able to recover from this error. Will add a catch to recover from the error, and then will work on solving the problem in the first place so that the error never occurs.
  • Mark: do tools to verify synchronization of the detector during a run exist?
  • Andrew: being worked on

DCS:

  • Athans:
    • showed demo code of sender <-> receiver applications for DCM DCS data.
    • See doc-dB 5497 for further details.
    • The next step is to add code to read back FEB status registers.
  • Chad:
    • dB access worked fine throughout the weekend.
    • working on writing data to the faults dB table.
    • will give a demonstration at the next meeting.

Shifters' report: all is well.